Blossom Potions, Inc., a pharmaceutical company, bought a machine at a cost of $2 million five years ago that produces pain-reliever medicine. The machine has been depreciated over the past five years, and the current book value is $832,000. The company decides to sell the machine now at its market price of $1 million. The marginal tax rate is 25 percent.

What are the relevant cash flows?

Relevant cash flow $



How do they change if the market price of the machine is $600,000 instead?

Relevant cash flow $
Answer #1

1: Cash flows =Selling price - Tax rate*(Selling price-Bookvalue)

= 1000000- 25%*(1000000-832000)

= 1000000-42000

= 958000

2:

Cash flows =Selling price - Tax rate*(Selling price-Bookvalue)

= 600000- 25%*(600000-832000)

=6000000+58000

= $ 658000